# Gamma family characterization and an alternative proof of Gini estimator unbiasedness

### v1

@@ 6 Concluding remarks
In this paper, we developed a general framework for studying the expectation of the Gini coefficient estimator through Laplace-transform techniques. This framework led naturally to a characterization of the gamma family within the class of nonnegative scale families, based on the invariance of exponentially tilted distributions under scaling.
As a direct application, we provided an alternative proof of the unbiasedness of the Gini estimator under gamma populations. Compared with existing proofs, the proposed approach is conceptually simpler and emphasizes structural properties, namely, scale invariance and stability under exponential tilting, rather than distribution-specific algebraic calculations.
Beyond the gamma case, the representation derived in Theorem 3.1 may prove useful for studying bias properties of Gini-type estimators under other distributional settings, including mixtures and generalized scale families. Future research could explore whether similar characterizations arise for other inequality measures or whether approximate unbiasedness results can be obtained for broader classes of distributions.
Overall, the results reinforce the central role of the gamma family in inequality measurement and provide further insight into the probabilistic foundations of Gini estimation.
@@ Acknowledgments
The research was supported in part by CNPq and CAPES grants from the Brazilian government.

### v2

@@ 6 Concluding remarks
In this paper, we derived a general representation for the expectation of the Gini coefficient estimator based on the Laplace transform of the underlying distribution and the Gini coefficient of its exponentially tilted version. This formulation yields a new characterization of the gamma family within the class of nonnegative scale families via a stability property under exponential tilting. As direct applications, we showed that the Gini estimator is biased for Poisson and geometric populations, while remaining unbiased for gamma models. By exploiting the derived bias expressions, we proposed plug-in bias-corrected estimators and assessed their finite-sample performance through Monte Carlo simulations, which confirmed substantial improvements over the uncorrected estimator.
@@ Acknowledgments
The research was supported in part by CNPq and CAPES grants from the Brazilian government. Helton Saulo gratefully acknowledges financial support from: (ii) the University of Brasília; and (ii) the Brazilian National Council for Scientific and Technological Development (CNPq), through Grant 304716/2023-5.